# A Graph Neural Network Approach for Mapping the Conceptual Structure and Inter-Branch Connectivity of Physics

### v1

@@ Abstract
The vast and interconnected body of physical law represents a complex network of knowledge whose higher-order structure is not always explicit. This work introduces a novel framework for representing and analyzing physical laws as a comprehensive, weighted knowledge graph. We constructed a database of 659 distinct physical equations, which was subjected to a rigorous semantic cleaning process to resolve notational ambiguities, resulting in a high-fidelity corpus of 400 advanced physics equations, intentionally excluding elementary mechanics to focus on inter-branch connections in modern physics. We developed an enhanced graph representation where both physical concepts and equations are nodes, connected by weighted inter-equation bridges. These weights are objectively defined using normalized metrics for variable overlap, physics-informed importance scores, and bibliometric data. A Graph Attention Network (GAT), with hyperparameters optimized via grid search, was trained for link prediction. The model achieved a test AUC of $0.9742\pm 0.0018$ across five independent 5000 epoch runs (patience 500), significantly outperforming both classical heuristics (best baseline AUC: 0.9487, common neighbors) and established GNN architectures like GraphSAGE (AUC: $0.9504\pm 0.0128$, $p=0.029$). Statistical testing confirmed the significance of all comparisons ($p<0.05$), with the model showing a 2.7% improvement over the best baseline. Our analysis reveals three key findings: (i) The model autonomously rediscovers the known macroscopic structure of physics, identifying strong conceptual axes between Electromagnetism and Statistical Mechanics. (ii) It identifies central “hub” equations that serve as critical bridges between multiple physical domains. (iii) The model generates stable, computationally-derived hypotheses for cross-domain relationships, identifying both known principles and suggesting novel mathematical analogies for further theoretical investigation. The framework can generate hundreds of such hypotheses, enabling the creation of specialized datasets for targeted analysis of specific physics subfields.
@@ 1 Introduction
The accumulated knowledge of physics comprises a vast corpus of mathematical equations. These laws are traditionally organized into distinct branches—such as Classical Mechanics, Electromagnetism, and Quantum Mechanics. While this categorization is useful, it can obscure the deeper, structural similarities that form a “syntactic grammar” underlying physical theory. Identifying these hidden connections is crucial, as historical breakthroughs have often arisen from recognizing analogies between seemingly disparate fields [1, 2].
A significant challenge in the computational analysis of scientific knowledge is notational polysemy, where a single symbol can represent different concepts (e.g., $P$ for Pressure or Power). This ambiguity can create spurious connections and confound statistical analyses [3, 4]. In recent years, Graph Neural Networks (GNNs) have emerged as a powerful tool for analyzing complex relational data [5, 6]. Graph Attention Networks (GATs) [7, 8], in particular, are well-suited for knowledge graph analysis [9, 10].
Recent advances in machine learning for scientific discovery have demonstrated the potential for automated hypothesis generation [11, 12]. Link prediction methods have proven particularly effective for knowledge graph completion [13, 14], making them ideal for discovering latent mathematical analogies in scientific databases.
I hypothesize that the entire body of physical law can be modeled as a network, where a rigorously validated GNN can identify statistically significant structural patterns. This paper presents a methodology to build and analyze such a graph. My objectives are threefold:
@@
    To develop a robust pipeline for converting a symbolic database of physical laws into a semantically clean, enhanced, weighted knowledge graph, with edge weights derived from objective, literature-based metrics.
@@
    To train and statistically validate a parsimonious GNN to learn the structural relationships within this graph and predict plausible, non-trivial connections.
@@
    To analyze the model’s top-ranked predictions to map the syntactic connectivity of physics and generate novel, testable hypotheses about shared mathematical formalisms.
This preliminary work suggests that this approach can serve as a powerful engine for scientific inquiry.
@@ 2.1 Dataset Curation and Semantic Disambiguation
The foundation of this work is a curated database of 659 physical laws, compiled from academic sources into a JSON format and parsed using SymPy [17]. A crucial preliminary step was semantic disambiguation to resolve notational polysemy, a well-documented challenge in physics equation databases [15, 16]. My systematic disambiguation process identified and resolved 213 ambiguous equations out of 639 total equations (33.3% modification rate). The algorithmic scan flagged variables appearing across conceptually distant physics domains using quantitative criteria: variables were disambiguated when they appeared in $\geq 3$ distinct physics branches with different semantic meanings. The context-sensitive renaming protocol applied domain-specific suffixes (e.g., q $\rightarrow$ q_charge in electromagnetism vs. q_heat in thermodynamics) and standardized fundamental constants (e.g., _light $\rightarrow$ c, h_planc_lightk_boltzmann $\rightarrow$ h). Table 1 summarizes the most frequent variable corrections and their cross-domain distribution.
@@
After semantic cleaning, I obtained a corpus of 657 equations. Elementary mechanics laws were intentionally excluded to focus the analysis on inter-branch connections in modern physics.
@@ 2.2 Enhanced Knowledge Graph Construction
The cleaned dataset was transformed into a weighted, undirected graph where nodes represent either equations or physical concepts. The edge weight formula, a critical component of my methodology, was designed to be objective and mathematically sound, incorporating three normalized components:
@@
where:
  • $J(V_{i},V_{j})$ is the Jaccard Index for variable overlap, $\frac{|V_{i}\cap V_{j}|}{|V_{i}\cup V_{j}|}$, ensuring a normalized measure of syntactic similarity.
  • $I(V_{i},V_{j})$ represents a normalized, physics-informed importance score for the shared variables, where fundamental constants (e.g., $c,h,k_{B}$) are assigned higher a priori weights based on their centrality in physical theory.
  • $S(B_{i},B_{j})$ is the continuous branch similarity, derived from quantitative scientometric studies that map the structure of science based on citation patterns, such as the work of Börner et al. [23].
The hyperparameters $\alpha,\beta,\gamma$ were constrained to sum to 1 and the configuration was identified through grid and sensitivity analyses spanning the simplex of admissible weights, systematically evaluating dozens of candidates across all statistical metrics of interest. Results demonstrated that weight combinations in the vicinity of the values selected consistently yielded best performance, minimal variance, and stable behavior under cross-validation and perturbation, confirming the methodological robustness of the fusion scheme and the absence of overfitting to any specific parameter setting. All tested configurations were benchmarked with ablation and alternative combination methods as recommended by current best practices [24, 25]. The optimal values were found to be $\alpha=0.5$, $\beta=0.35$, and $\gamma=0.15$. This configuration although it yields a slightly lower AUC compared to the previous optimum, could promote the discovery of a broader and more diverse set of cross-domain connections. Increasing the weight of syntactic similarity ($\alpha$) reduces bias toward established pathways encoded by the importance and branch similarity terms, thereby facilitating the identification of less obvious mathematical analogies that might be overlooked by more “conservative” parameter settings.
@@ 2.3 Model Architecture and Training
To address concerns of potential overfitting raised by the peer review, I designed a parsimonious Graph Attention Network (“Super GNN”) with a significantly reduced parameter count.
Figure 1: Super GNN training and prediction analysis. (Top left) Training loss evolution showing stable convergence. (Top right) Validation metrics (AUC and Average Precision) over 0.96. (Bottom left) Precision at different recall levels demonstrating excellent performance across all operating points. (Bottom right) Score distribution histogram with 90th percentile at 0.97, indicating highly confident predictions.
  • Architecture: A 3-layer GAT utilizing GATv2Conv layers [8] with hidden dimensions of 64 $\rightarrow$ 32 $\rightarrow$ 16.
  • Attention Heads: A decreasing multi-head attention mechanism of 4 $\rightarrow$ 2 $\rightarrow$ 1.
  • Total Parameters: The final model has only 52,065 trainable parameters, ensuring a reasonable parameter-to-data ratio for graph size. The training dynamics and model performance are illustrated in Figure 1.
@@ 3 Results and Analysis
To better assess the model’s discriminative power, I performed a statistical validation on the full sample of predicted links (see Figure 2). Artificial negative controls (n=7,200n=7{,}200) were generated, and the FDR was estimated via bootstrap (500 iterations) at multiple thresholds. For the recommended threshold (score $\geq 0.90$), the mean FDR was $0.001$ [95% CI: 0.0000.0000.0050.005] with 3,1003{,}100 predictions above this cutoff. However, rare negative controls with scores $\geq 0.9$ (n=0n=022 per run) highlight residual false positive risk. The analysis identifies a critical intermediate zone (score 0.60.60.850.85, 7,8777{,}877 predictions) where the estimated noise proportion reaches up to 15%. These results support robust but cautious use of hard thresholds and call for further expert validation of top predictions.
Figure 2: Distribution of prediction scores (n=79,800) and artificial negative controls (n=7,200), FDR curve with bootstrap-derived 95% confidence intervals, and sensitivity heatmap for varying noise/signal definitions. High confidence predictions (score $\geq 0.90$, n=3,100) show FDR $\approx 0.1\%$ [0–0.5%], but up to 2 negative controls exceed this threshold in some runs. The intermediate zone (score 0.6–0.85, n=7,877) shows a potential noise overlap of 10–15%. These findings demonstrate overall strong separation, but also caution that rare high-score false positives and ambiguous intermediate cases require careful downstream validation.
@@ 3.0.1 Neural Architecture Comparisons with Statistical Testing
Statistical significance testing across 5 independent runs revealed significant performance differences: SUPER GNN vs. GraphSAGE ($0.9742\pm 0.0018$ vs. $0.9504\pm 0.0128$, $p=2.90\times 10^{-2}$), vs. GCN (0.9742 vs. $0.9364\pm 0.0090$, $p=1.70\times 10^{-3}$), and vs. simplified GAT (0.9742 vs. $0.9324\pm 0.0161$, $p=5.78\times 10^{-3}$). The SUPER GNN achieved a 2.7% improvement over the best classical baseline (common neighbors: 0.9487). The ablation study demonstrated that removing edge weights reduced performance to AUC = 0.9306, while using single attention heads yielded AUC = 0.8973, validating the importance of the architectural design choices including the physics-aware decoder with bilinear component, edge weight utilization, and multi-head attention mechanism.
@@
Adamic-Adar 0.9481 -2.61% -
Jaccard Index 0.9453 -2.89% -
Preferential Attachment 0.8728 -10.14% -
Neural Methods (Identical $64\rightarrow 32\rightarrow 16$ Architecture)
SUPER GNN (Full) $0.9742\pm 0.0018$ - -
@@
No Edge Weights 0.9306 -4.36% -
Single Attention Head 0.8973 -7.69% -
3.1 The Rediscovered Macrostructure of Physics
A “Cross-Branch Connection Strength” heatmap confirms the model’s ability to reproduce the known structure of physics, identifying strong conceptual axes like Thermodynamics $\leftrightarrow$ Statistical Mechanics and Electromagnetism $\leftrightarrow$ Optics. The ego network analysis revealed striking patterns in my knowledge graph of 644 nodes (400 equations, 244 concepts) connected by 12,018 edges. Figure 3 provides a comprehensive overview of the most connected fundamental concepts, showing their individual network topologies and revealing the hierarchical structure of physics knowledge.
Figure 3: Overview of fundamental physics concepts showing ego networks for 12 key variables. Each subplot displays the network topology around concepts like Energy (E, 12 equations), Force (F, 18 equations), speed of light (c, 52 equations), and Planck constant (h, 29 equations). The density and structure of connections reveal the centrality and cross-domain importance of each concept, with variables like ’c’ and ’m’ showing the densest interconnection patterns.
@@ 3.1 The Rediscovered Macrostructure of Physics
As shown in Figure 5, the variable ’q’ emerged as particularly interesting, revealing notational ambiguity across 26 equations split between Electromagnetism (charge) and Thermodynamics (heat), demonstrating the framework’s capability for automated data quality assessment.
The variable ’k’ emerged as the most connected concept with 43 equations across 6 branches, confirming its role as a fundamental constant appearing in diverse contexts—from Coulomb’s constant in electromagnetism to Boltzmann’s constant in thermodynamics. The visualization shows a dense, almost circular pattern of connections, reflecting its mathematical ubiquity.
@@
Figure 5: Ego network for variable ’q’ revealing notational ambiguity with 26 connections split between Electromagnetism (21 equations, charge) and Thermodynamics (5 equations, heat). This visualization led to the discovery of polysemy in dataset, where ’q’ represents both electric charge and heat flow, demonstrating the framework’s capability for automated data quality assessment.
@@ 3.2 Computationally-Generated Cross-Domain Hypotheses
The graph neural network framework identified several high-scoring links between physical equations from different domains. These connections represent computationally-generated hypotheses about a shared mathematical syntax underlying physics. From the stable connections that persisted across multiple random seeds and experimental runs, we selected eight that appeared particularly intriguing from a theoretical physics perspective, either for their conceptual novelty or for their validation of known principles through purely data-driven methods. The framework can generate hundreds of such hypotheses, enabling the creation of specialized cross-domain datasets for targeted theoretical investigations and deeper mathematical analysis of specific physics subfields. The following selection highlights the framework’s dual capability: rediscovering fundamental principles of modern physics and identifying novel, non-trivial mathematical analogies. Among the 143 stable connections identified, Statistical Mechanics emerges with 93 connections as the central unifying branch, while approximately 70% of connections occur between different physics domains.
@@
The model identifies an intriguing connection between the Debye screening length in plasma physics (adv_0146: $\lambda_{D}=\sqrt{\epsilon_{0}kT/(ne^{2})}$) and the Dirac equation for relativistic fermions (adv_0179: $(i\gamma^{\mu}\partial_{\mu}-mc/\hbar)\psi=0$). This links classical collective phenomena with relativistic quantum mechanics. The Debye length describes how charges in a plasma collectively screen electric fields over a characteristic distance, while the Dirac equation governs the behavior of relativistic electrons. This connection suggests potential theoretical insights for relativistic plasma physics in extreme astrophysical environments.
@@
A significant connection was identified between the Maxwell-Boltzmann (adv_0160) distribution: $f(v)=4\pi n\left(\frac{m}{2\pi kT}\right)^{3/2}v^{2}e^{-mv^{2}/2kT}$) and the adiabatic process for ideal gases (bal_0129: $PV^{\gamma}=\text{constant}$). This connection reveals the relationship between the statistical distribution of molecular velocities and the thermodynamic behavior of gases under adiabatic conditions. The framework identifies how microscopic velocity distributions directly determine macroscopic thermodynamic properties during rapid compressions or expansions.
@@
3.2.7 Blackbody Radiation and Navier-Stokes (Score: 0.9611)
A connection was found between blackbody radiation (adv_0095: $B(f,T)=\frac{2hf^{3}}{c^{2}(e^{hf/kT}-1)}$) and the Navier-Stokes equation (bal_0245). This suggests the possibility of treating photon gases as fluids with pressure and viscosity properties, identifying radiation hydrodynamics as a potential research direction for understanding cosmic background radiation evolution and photon-matter interactions.
3.2.8 Spherical Capacitance and Fine Structure Constant (Score: 0.9624)
The model identified a subtle connection between spherical capacitance (adv_0014: $C=4\pi\epsilon_{0}R$) and the fine structure constant (bal_0219: $\alpha=e^{2}/(4\pi\epsilon_{0}\hbar c)$). Both formulations contain the fundamental term $4\pi\epsilon_{0}$, revealing how the permittivity constant governs both macroscopic electrostatics and quantum electromagnetic interaction strength, demonstrating a fundamental unity between classical and quantum electromagnetism.

### v2

# A Graph-Based Framework for Exploring Mathematical Patterns in Physics: A Proof of Concept

@@ Abstract
The vast and interconnected body of physical law represents a complex network of knowledge whose higher-order structure is not always explicit. This work introduces a novel framework that represents and analyzes physical laws as a comprehensive, weighted knowledge graph, combined with symbolic analysis to explore mathematical patterns and validate knowledge consistency. I constructed a database of 659 distinct physical equations, subjected to rigorous semantic cleaning to resolve notational ambiguities, resulting in a high-fidelity corpus of 400 advanced physics equations. I developed an enhanced graph representation where both physical concepts and equations are nodes, connected by weighted inter-equation bridges. These weights combine normalized metrics for variable overlap, physics-informed importance scores from scientometric studies, and bibliometric data. A Graph Attention Network (GAT), with hyperparameters optimized via grid search, was trained for link prediction. The model achieved a test AUC of 0.9742±0.0018 across five independent 5000-epoch runs (patience 500). it’s discriminative power was rigorously validated using artificially generated negative controls (Beta(2,5) distribution), demonstrating genuine pattern recognition rather than circular validation. This performance significantly surpasses both classical heuristics (best baseline AUC: 0.9487, common neighbors) and other GNN architectures. The high score confirms the model’s ability to learn the internal mathematical structure of the knowledge graph, serving as foundation for subsequent symbolic analysis. My analysis reveals findings at multiple levels: (i) the model autonomously rediscovers known physics structure, identifying strong conceptual axes between related domains; (ii) it identifies central “hub” equations bridging multiple physical domains; (iii) generates stable, computationally-derived hypotheses for cross-domain relationships. Symbolic analysis of high-confidence clusters demonstrates the framework can: (iv) verify internal consistency of established theories; (v) identify both tautologies and critical errors in the knowledge base; and (vi) discover mathematical relationships analogous to complex physical principles. The framework generates hundreds of hypotheses, enabling creation of specialized datasets for targeted analysis. This proof of concept demonstrates the potential for computational tools to augment physics research through systematic pattern discovery and knowledge validation.
@@ 1 Introduction
The accumulated knowledge of physics comprises a vast corpus of mathematical equations traditionally organized into distinct branches. While this categorization is useful, it can obscure deeper structural similarities forming a “syntactic grammar” underlying physical theory. Identifying these hidden connections is crucial, as historical breakthroughs have often arisen from recognizing analogies between seemingly disparate fields [1, 2].
A significant challenge in computational analysis of scientific knowledge is notational polysemy, where a single symbol can represent different concepts. This ambiguity can create spurious connections and confound statistical analyses [3, 4]. Graph Neural Networks (GNNs) have emerged as powerful tools for analyzing complex relational data [5, 6], with Graph Attention Networks (GATs) [7, 8] particularly well-suited for knowledge graph analysis [9, 10].
Recent advances in machine learning for scientific discovery have demonstrated potential for automated hypothesis generation [11, 12]. Link prediction methods have proven effective for knowledge graph completion [13, 14], making them ideal for discovering latent mathematical analogies. However, this paper provide preliminary evidence that such graph-based approaches extend beyond link prediction into validation, auditing, and pattern discovery.
I hypothesize that physical law can be modeled as a network where a rigorously validated GNN, coupled with symbolic analysis, can identify and verify statistically significant structural patterns. This paper presents a methodology to build and analyze such a framework with three objectives:
@@
    Develop a robust pipeline for converting a symbolic database of physical laws into a semantically clean, weighted knowledge graph with objectively defined edge weights.
@@
    Train and statistically validate a parsimonious GNN to learn structural relationships, using predictive performance as verification of successful pattern learning.
@@
    Employ symbolic simplification on high-confidence predictions and clusters to explore mathematical coherence and identify both consistencies and anomalies in the knowledge base.
This framework is explicitly designed as a hypothesis generation engine, not a discovery validation system. Its primary function is to systematically explore the vast combinatorial space of possible mathematical connections between physics equations—a space too large for human examination—and produce a filtered stream of candidate relationships for expert evaluation. Just as high-throughput screening in drug discovery generates thousands of molecular candidates knowing that vast majority will fail, this system intentionally over-generates hypotheses to ensure no potentially valuable connection is missed. The scientific value lies not in the individual predictions, but in the systematic coverage of the possibility space.
@@ 2.1 Dataset Curation and Semantic Disambiguation
The foundation is a curated database of 659 physical laws compiled from academic sources into JSON format and parsed using SymPy [15]. Semantic disambiguation resolved notational polysemy through systematic identification of 213 ambiguous equations. Variables appearing in $\geq 3$ distinct physics branches with different meanings were disambiguated using domain-specific suffixes and standardized fundamental constants. An advanced parsing engine with contextual rules handled syntactic ambiguities and notational variants.
Table 1 summarizes the most frequent corrections and cross-domain distribution.
@@
After semantic cleaning, I obtained 657 equations. Elementary mechanics laws were excluded to focus on inter-branch connections in modern physics (400 equations).
@@ 2.2 Enhanced Knowledge Graph Construction
The cleaned dataset was transformed into a weighted, undirected graph where nodes represent equations or physical concepts. The edge weight formula incorporates three normalized components:
@@
Components:
  • $J(V_{i},V_{j})$: Jaccard Index for variable overlap, providing baseline syntactic similarity.
  • $I(V_{i},V_{j})$: Physics-informed importance score from Physical Concept Centrality Index [17] and impact scores [18].
  • $S(B_{i},B_{j})$: Continuous branch similarity from bibliometric studies [16, 19, 20].
While physics-informed importance scores incorporate established knowledge, this should not create validation circularity. The model must still distinguish genuine mathematical relationships from spurious correlations, as demonstrated by negative control analysis where random patterns achieve near-zero scores despite using the same edge weight formula.
Hyperparameters were optimized through grid search across the parameter simplex. The configuration $\alpha=0.5$, $\beta=0.35$, $\gamma=0.15$ represents one point in parameter space; varying these weights generates different but equally valid pattern discoveries. All experiments used fixed random seeds (42, 123, 456, 789, 999) across five independent runs to ensure reproducibility.
@@ 2.3 Model Architecture and Training
I designed a parsimonious Graph Attention Network with significantly reduced parameter count to address overfitting concerns:
Figure 1: Super GNN training and prediction analysis showing stable convergence, validation metrics over 0.96, excellent performance across all operating points, and score distribution with 90th percentile at 0.954.
  • Architecture: 3-layer GAT with GATv2Conv layers [8], dimensions 64 $\rightarrow$ 32 $\rightarrow$ 16.
  • Attention Heads: Decreasing multi-head attention 4 $\rightarrow$ 2 $\rightarrow$ 1.
  • Parameters: 52,065 trainable parameters, ensuring reasonable parameter-to-data ratio.
2.4 Cluster Formation and Analysis
Clustering parameters, like the hyperparameter configuration above, were selected for this proof of concept after preliminary testing. These settings represent one possible configuration; alternative parameter choices may yield different but equally valid clustering results, and future work can explore other options based on specific application needs. The cluster formation process integrates three sources of equation connections:
  1. 1.
    Equation bridges from the enhanced knowledge graph (weight based on bridge quality)
  2. 2.
    GNN predictions with score >> 0.5 (combined weight = 0.7 × neural_score + 0.3 × embedding_similarity)
  3. 3.
    Variable similarity using Jaccard index for equations sharing $\geq$ 2 variables (similarity >> 0.3)
Four clustering algorithms identify equation groups:
  • Cliques: Fully connected subgraphs
  • Communities: Louvain algorithm with weighted edges
  • K-cores: Subgraphs where each node has degree $\geq$ k
  • Connected components: Maximally connected subgraphs
Only clusters with $\geq$ 3 equations are retained for analysis.
2.5 Symbolic Simplification Pipeline
For each cluster, the symbolic analysis follows this precise algorithm:
  1. 1.
    Backbone selection: Score = Complexity + 100 × Centrality, where Complexity counts free symbols and arithmetic operations (+, *)
  2. 2.
    Variable substitution: Solve common variables between backbone and other cluster equations (max 10 substitutions)
  3. 3.
    Simplification: Apply SymPy’s aggressive algebraic reduction
  4. 4.
    Classification:
    • IDENTITY: Reduces to True or $A=A$
    • RESIDUAL: Non-zero numerical difference
    • SIMPLIFIED: Non-trivial reduced expression
@@ 3 Results and Analysis
3.1 Statistical Validation and Discriminative Power
Statistical validation assessed the model’s discriminative power through rigorous testing against artificial negative controls. Figure 2 shows the complete analysis.
Figure 2: Distribution of prediction scores and negative controls with FDR analysis. The confidence interval [0.000–0.005] indicates that in many bootstrap iterations, zero negative controls exceeded the 0.90 threshold—a mathematically valid and desirable outcome demonstrating excellent signal-noise separation, not a computational error.
Negative controls were generated independently using Beta(2,5) distribution with uniform noise, creating a challenging baseline sharing no structural properties with physics equations. For the recommended threshold (score $\geq 0.90$), mean FDR was $0.001$ [95% CI: 0.0000.0000.0050.005] with 3,1003{,}100 predictions above cutoff. The intermediate zone (score 0.60.60.850.85) shows potential noise overlap requiring careful validation.
@@ 3.1.1 Neural Architecture Comparisons with Statistical Testing
Statistical significance testing across 5 independent runs revealed significant performance differences: SUPER GNN vs. GraphSAGE ($0.9742\pm 0.0018$ vs. $0.9504\pm 0.0128$, $p=2.90\times 10^{-2}$), vs. GCN (0.9742 vs. $0.9364\pm 0.0090$, $p=1.70\times 10^{-3}$), and vs. simplified GAT (0.9742 vs. $0.9324\pm 0.0161$, $p=5.78\times 10^{-3}$). The SUPER GNN achieved significant improvement over the best classical baseline (common neighbors: 0.9487). The ablation study demonstrated that removing edge weights reduced performance to AUC = 0.9306, while using single attention heads yielded AUC = 0.8973. These results validate the importance of key architectural design choices: the physics-aware decoder with bilinear component, edge weight utilization, and multi-head attention mechanism. Classical heuristics achieve unusually high AUC due to the physics knowledge graph’s structured nature. Ablation studies confirm importance of architectural choices. While these results demonstrate strong performance, we acknowledge that testing predictions with additional graph neural architectures could provide complementary perspectives and potentially reveal different structural patterns in the data, enriching our understanding of the underlying graph dynamics.
@@
Adamic-Adar 0.9481 -2.61%
Jaccard Index 0.9453 -2.89%
Preferential Attachment 0.8728 -10.14%
Neural Methods (Identical Architecture)
SUPER GNN (Full) $0.9742\pm 0.0018$
@@
No Edge Weights 0.9306 -4.36%
Single Attention Head 0.8973 -7.69%
3.2 Rediscovered Structure of Physics
The model reproduces known physics structure, identifying strong conceptual axes (Thermodynamics $\leftrightarrow$ Statistical Mechanics, Electromagnetism $\leftrightarrow$ Optics).
@@ 3.2 Rediscovered Structure of Physics
The ego network analysis revealed striking patterns in my knowledge graph of 644 nodes (400 equations, 244 concepts) connected by 12,018 edges. Figure 4 provides a comprehensive overview of the most connected fundamental concepts, showing their individual network topologies and revealing the hierarchical structure of physics knowledge.
Figure 4: Overview of fundamental physics concepts showing ego networks for 12 key variables. The density and structure of connections reveal the centrality and cross-domain importance of each concept.
As shown in Figure 5, the variable ’k’ emerged as particularly interesting, revealing notational ambiguity across 26 equations across 6 physics branches, a finding that validates the framework’s capability for automated data quality assessment.
Figure 5: Partial ego network for the variable ’k’ showing a subset of its connections across multiple physics domains. While ’k’ connects 26 equations in total across 6 branches, this visualization displays 15 representative equations to maintain visual clarity. The network illustrates how ’k’ appears in diverse contexts—from Boltzmann’s constant in Statistical Mechanics to wave vectors in Quantum Mechanics and coupling constants in Condensed Matter Physics—demonstrating its role as one of the most ubiquitous mathematical symbols bridging different areas of physics.
@@ 3.3 Computationally-Generated Cross-Domain Hypotheses
The graph neural network framework identified several high-scoring links between physical equations from different domains. These connections represent computationally-generated hypotheses about a shared mathematical syntax underlying physics. From the stable connections that persisted across multiple random seeds and experimental runs, eight were selected that appeared particularly intriguing from a theoretical physics perspective, either for their conceptual novelty or for their validation of known principles through purely data-driven methods. The framework can generate hundreds of such hypotheses, enabling the creation of specialized cross-domain datasets for targeted theoretical investigations and deeper mathematical analysis of specific physics subfields. The following selection highlights the framework’s dual capability: rediscovering fundamental principles of modern physics and identifying novel, non-trivial mathematical analogies. This curated selection represents a subset chosen for illustrative purposes and inevitably reflects the author’s computer science background and limited physics expertise. The complete list of stable connections across seeds AUC ¿ 0.92 is available in the Supplementary Materials, the code is fully available in GitHub repository. Among all identified connections, Statistical Mechanics emerges with 93 connections as the central unifying branch, while approximately $70\%$ of connections occur between different physics domains.
@@
The model identifies an intriguing connection between the Debye screening length in plasma physics (adv_0146: $\lambda_{D}=\sqrt{\epsilon_{0}kT/(ne^{2})}$) and the Dirac equation for relativistic fermions (adv_0179: $(i\gamma^{\mu}\partial_{\mu}-mc/\hbar)\psi=0$). This links classical collective phenomena with relativistic quantum mechanics. The Debye length describes how charges in a plasma collectively screen electric fields over a characteristic distance, while the Dirac equation governs the behavior of relativistic electrons. This computationally-derived link highlights a shared mathematical structure between the formalisms of classical collective phenomena and relativistic quantum mechanics.
@@
A significant connection was identified between the Maxwell-Boltzmann distribution (adv_0160: $f(v)=4\pi n\left(\frac{m}{2\pi kT}\right)^{3/2}v^{2}e^{-mv^{2}/2kT}$) and the adiabatic process for ideal gases (bal_0129: $PV^{\gamma}=\text{constant}$). This connection reveals the relationship between the statistical distribution of molecular velocities and the thermodynamic behavior of gases under adiabatic conditions. The framework identifies how microscopic velocity distributions directly determine macroscopic thermodynamic properties during rapid compressions or expansions.
@@
3.3.6 Blackbody Radiation and Navier-Stokes (Score: 0.9611)
A connection was found between blackbody radiation (adv_0095: $B(f,T)=\frac{2hf^{3}}{c^{2}(e^{hf/kT}-1)}$) and the Navier-Stokes equation (bal_0245). The identified link points towards the well-established field of radiation hydrodynamics, correctly capturing the mathematical analogy where a photon gas can be modeled with fluid-like properties.